# Title of the Paper

### v2

@@ Abstract
This paper studies how payoff heterogeneity affects the value of information in beauty contest games. I show that public information is detrimental to welfare if and only if agents’ Katz-Bonacich centralities exhibit specific forms of heterogeneity, stemming from the network of coordination motives. A key insight is that agents may value the commonality of information so differently that some are harmed by their neighbors knowing what others know. Leveraging this insight, I also show that when the commonality of information is endogenously determined through information sharing, the equilibrium degree of information sharing can be inefficiently low, even without sharing costs.
@@ 1 Introduction
Many socioeconomic interactions are shaped simultaneously by incomplete information and complex strategic interdependencies. A user deciding on technology adoption often faces uncertainty about its intrinsic value and payoffs contingent on adoption patterns within their social network. An investor formulates investment strategies based on unknown economic fundamentals while anticipating the behavior of other market participants, especially influential ones. Within organizations, a worker facing task uncertainty often needs to coordinate their actions with specific individuals or teams, as dictated by formal and informal structures.
Despite the recognized prevalence of information and payoff heterogeneity, their interplay remains relatively underexplored. This paper bridges this gap by examining how payoff heterogeneity arising from networked interactions affects the welfare implications of public information. For this purpose, I introduce a beauty contest game incorporating a flexible form of coordination motives represented by a weighted and directed network. Agents incur quadratic losses from deviations of their actions from a common payoff state and the actions of their specific neighbors. Agents do not directly observe the state but receive private and public signals.
The main result of this paper provides a geometric characterization of interaction networks under which public information provision can be detrimental to welfare. This result might seem counterintuitive, as public information is typically beneficial because it informs an agent about the payoff state and what her neighbors know. However, public information also informs her neighbors about what the neighbors of her neighbors know. This increased commonality of information can distort her neighbors’ actions away from the state, thereby imposing a negative externality on her. Depending on the network topology, such negative externalities can dominate the informational gains from public information.
While the finding that public information can harm welfare in beauty contest games resonates with the seminal work of Morris and Shin, (2002), the underlying economic mechanism identified here is distinct. The anti-transparency result in Morris and Shin hinges on a specific welfare criterion independent of coordination losses; only deviations from the state affect aggregate welfare. The result in this paper requires no such assumption on the welfare function; instead, the detrimental effect of public information necessitates a sufficient degree of heterogeneity in coordination motives.
The formal analysis employs a linear-quadratic-Gaussian framework. Linear best responses, combined with Gaussian signals, ensure the existence of a unique linear equilibrium. In this equilibrium, the weight placed on the public signal is proportional to an agent’s Katz-Bonacich centrality, which captures her direct and indirect coordination motives (i.e., coordination motives of neighbors, neighbors’ neighbors, etc.). Thus, more central agents are more responsive to the public signal than less central ones in equilibrium.
The first result shows that public information provision decreases an individual agent’s equilibrium payoff if and only if the agent’s Katz-Bonacich centrality is sufficiently low relative to an aggregate measure of her neighbors’ centralities. For example, in core-periphery networks, harm from public information typically arises for peripheral agents who intend to coordinate with agents in a highly cohesive core. The core agents place substantial weight on public information due to their strong coordination motives, sometimes at the expense of aligning their actions with the payoff state. [^1] This observation is reminiscent of the anti-transparency result of Morris and Shin, (2002). This deviation by the core agents, in turn, makes it more difficult for the peripheral agents to simultaneously align their actions with the state and the core agents’ actions.
Public information may also decrease aggregate welfare. These detrimental welfare effects tend to arise in networks characterized by a few highly central agents serving as coordination hubs for many less-central agents. In core-periphery networks with a small, cohesive core and a large periphery, public information benefits core agents while potentially harming peripheral agents. Aggregate welfare declines if the losses incurred by the peripheral agents exceed the gains accruing to the core agents.
Finally, applying insights from the main analysis, I find a novel source of inefficiency in endogenous information sharing. Consider an extension where an agent initially possesses a signal with the potential for public disclosure and can commit to sharing it publicly. In this setting, the agent may choose not to share the signal in equilibrium even when its public disclosure would be socially desirable. In core-periphery networks, a peripheral agent may prefer to withhold her signal to prevent core agents from being overly responsive should the signal be shared. Such a decision by the peripheral agent can lead to substantial welfare losses, as the potential benefits to core agents are forgone.
@@ 2 Literature
This paper contributes to the literature on the social value of information. In their seminal work, Morris and Shin, (2002) developed the beauty contest model and showed that public information can harm welfare. Their model has been applied to diverse contexts, including technological investment (Angeletos and Pavan,, 2004), asset pricing (Allen et al.,, 2006), and political leadership (Dewan and Myatt,, 2008). Angeletos and Pavan, (2007) and Ui and Yoshizawa, (2015) characterized the social value of public information in more general linear-quadratic games. Other notable extensions include work by Colombo et al., (2014) on endogenous acquisition of private information, Bergemann and Morris, (2013) on optimal Bayes correlated equilibria, and Huo and Pedroni, (2020) on dynamic information. Pavan and Vives, (2015) and Angeletos and Lian, (2016) provide an excellent overview of this literature.
Almost all existing work in this literature assumes no payoff heterogeneity. [^2] There are a few exceptions. The solution method for dynamic beauty contest games developed by Huo and Pedroni, (2020) applies to network games. Recent work by Miyashita and Ui, (2023) proposed a semi-definite programming approach to information design problems in network games. Also, Morris and Shin’s anti-transparency result hinges on specific assumptions about the welfare function, particularly its independence from coordination losses (Angeletos and Pavan,, 2004; Hellwig,, 2005; Woodford,, 2005). In contrast, I study environments where agents have heterogeneous coordination motives and the welfare measure incorporates coordination losses. The detrimental effect of public information identified in this paper arises from a novel mechanism that necessitates payoff heterogeneity.
The paper also contributes to the literature on endogenous information acquisition within the beauty contest framework. Hellwig and Veldkamp, (2009) and Myatt and Wallace, (2012) developed flexible yet tractable information acquisition models where information publicity and precision are endogenously determined. More recently, Herskovic and Ramos, (2020) studied information acquisition from neighbors in endogenously formed networks.
A key feature of these models is that “agents want to know what others know” due to their coordination motives; specifically, the marginal benefit of observing a signal increases with the number of other observers. This is not necessarily true when coordination motives are heterogeneous, as agents may not want specific other agents to observe their signal. Using a simple model of endogenous information sharing, I show that such non-monotonicity can cause inefficiency in the equilibrium degree of information sharing.
A few recent papers have studied endogenous information acquisition in network games. Myatt and Wallace, (2019) introduced heterogeneous coordination motives to Myatt and Wallace, (2012). Leister, (2020) examined endogenous acquisition of private information in the network game of Ballester et al., (2006). [^3] In related models, Calvó-Armengol and de Martí, (2009), Calvó-Armengol et al., (2015), and Myatt and Wallace, (2018) also consider endogenous information acquisition with payoff heterogeneity. Their main focus is to characterize how equilibrium information acquisition depends on the network of payoff interdependencies, which is complementary to but different from mine. [^4] For instance, Proposition 2 of Myatt and Wallace, (2019) provides an equilibrium characterization similar to one in this paper: Agents with higher Katz-Bonacich centrality are more responsive than those with lower centrality to relatively public signals (whose publicity is endogenous in their model).
Finally, in a companion paper (Murayama,, 2025), I generalized the anti-transparency result of Morris and Shin, (2002) to large network games. That paper shows that when welfare solely depends on the action-state distance, public information provision harms welfare if and only if the mean square of the Katz-Bonacich centrality is sufficiently large. Interestingly, these detrimental effects may arise even when agents’ coordination motives are positive but arbitrarily small on average, highlighting the importance of indirect coordination motives.
@@ 3 Model
I introduce a beauty contest game in which agents’ coordination motives are represented as a weighted and directed network.
@@
Information Structure. The information structure is standard, following Morris and Shin, (2002). The payoff state $\theta$ follows an improper uniform prior distribution over the real line. Each agent $i$ observes a signal vector $\mathbf{s}_{i}=(x_{i},y)$, where $x_{i}$ is a private signal observed only by $i$, and $y$ is a public signal observed by all agents. Specifically, I define $x_{i}\equiv\theta+\varepsilon_{i}$ and $y\equiv\theta+\varepsilon_{0},$ where $\varepsilon_{i}$ and $\varepsilon_{0}$ are normally distributed with
$\mathbb{E}[\varepsilon_{i}]=\mathbb{E}[\varepsilon_{0}]=0,\mathrm{Var}[\varepsilon_{i}]=\tau_{x}^{-1},\mathrm{Var}[\varepsilon_{0}]=\tau_{y}^{-1},$
and the noise terms and $\theta$ are all mutually independent. The relative precision of private signals is denoted by $\gamma\equiv\tau_{x}/(\tau_{x}+\tau_{y})\in(0,1)$.
@@
where $\mathbb{E}_{i}[\cdot]\equiv\mathbb{E}[\cdot\mid\mathbf{s}_{i}]$ is the conditional expectation operator, conditional on $i$’s information $\mathbf{s}_{i}$. Consider linear strategies of the form $a_{i}=b_{i}y+(1-b_{i})x_{i}$ for all $i\in N$ [^5] Focusing on this form of linear strategies is without loss of generality. In Section 6.1, I show that the equilibrium slopes must sum to one for all agents in any linear equilibrium. Plugging these linear strategies into (2) yields
@@
[…]
@@
In equilibrium, agent $i$’s direct and indirect coordination motives are positively associated with her responsiveness to the public signal. Intuitively, when $i$’s coordination motive strengthens, the commonality of the public signal becomes more valuable for aligning her action with her neighbors’ actions. Moreover, if $i$’s neighbors have stronger coordination motives, they will respond more to the public signal, further increasing the value of the public signal as a coordination device. Thus, not only an agent’s own coordination motive but also those of her neighbors (and neighbors’ neighbors, etc.) positively affect her responsiveness to the public signal. [^7] The relative precision of private information, $\gamma$, acts as the equilibrium discount factor in the Katz-Bonacich centrality. This role arises from how different orders of agents’ expectations about $\theta$ (e.g., $\mathbb{E}_{i}[\theta]$ and $\mathbb{E}_{i}[\mathbb{E}_{j}[\theta]])$ respond to changes in $\gamma$. A more detailed intuition is provided in Murayama, (2025).
@@ 4 Results
The main results show that public information provision can be detrimental to an individual agent and aggregate welfare, depending on the geometry of interaction networks. Most proofs and derivations are provided in Section 6.
@@ 4.1 Individual Payoff
Let $\Delta U_{i}(\tau_{x},\tau_{y},G)\equiv U_{i}(\tau_{x},\tau_{y},G)-U_{i}(\tau_{x},0,G)$ be the effect of public information provision on agent $i$ [^9] Abusing notation, I write $U_{i}(\tau_{x},0,G)=-\tau_{x}^{-1}(1+d_{i}^{out})$ for agent $i$’s equilibrium payoff in the absence of public information. The following result is immediate from (11).
@@
The detrimental effect on Ann, stemming from the increased commonality of a signal between Bob and Carol, is reminiscent of a mechanism highlighted by Morris and Shin, (2002). Morris and Shin show that when agents’ common coordination motive (analogous to $\beta$ in my model) is strong enough and the relative precision of private information is sufficiently high, public information provision increases the action-state distance, i.e., $\mathbb{E}[(a^{*}_{i}-\theta)^{2}]$. Similarly, the increased commonality of information between Bob and Carol can make Bob’s equilibrium action distant from $\theta$ [^11] Intuitively, this occurs because Bob and Carol’s coordination motive is so large that they choose to be much more responsive to the public signal than they would be when their objectives are to align their actions only with $\theta$. This is not desirable for Ann, whose objective is to simultaneously align her action with $\theta$ and Bob’s action.

### v3

# Network Heterogeneity and Value of Information

@@ Abstract
Does greater connectivity enhance the value of public information? I study a networked beauty contest game where agents balance adaptation to the fundamental with local coordination. The analysis reveals a stark non-monotonicity: while public disclosure improves welfare when interactions are uniform, regardless of their intensity, it can be detrimental in core-periphery structures. This welfare loss stems from a distortion driven by the core, where core agents over-respond to a noisy public signal, forcing peripheral neighbors to absorb this volatility to maintain alignment. These findings suggest that standard transparency policies can backfire in tiered markets where dominant hubs propagate excess volatility.
@@ 1 Introduction
Many economic interactions exhibit local strategic complementarities, where agents seek to coordinate with a specific subset of partners rather than with the entire population. In production networks, for instance, a firm’s optimal pricing depends on the pricing decisions of its direct suppliers and distributors. Similarly, in over-the-counter (OTC) financial markets, dealers’ willingness to hold inventory depends on their ability to rebalance positions through their specific inter-dealer connections.
While coordination motives are defined by a local neighborhood, agents’ beliefs are often anchored by public information. Central bank inflation forecasts shape firms’ macroeconomic expectations, which in turn inform their pricing decisions (Coibion et al., 2018, 2020). In OTC markets, regulators utilize public stress tests to stabilize lending, providing a system-wide signal that reduces opacity and coordinates market beliefs (Petrella and Resti, 2013; Morgan et al., 2014; Flannery et al., 2017).
Standard intuition suggests that public disclosures are most effective when coordination incentives are uniformly strong, as public signals serve as a focal point for global alignment. Where interactions are weak or sparse, the social value of public information is therefore expected to diminish. This paper revisits this intuition by asking: How does the shape of interaction networks alter the welfare implications of public information?
To address this question, I study a beauty contest game embedded in a general interaction network. Agents face two competing objectives: an adaptation motive to match the fundamental state and a coordination motive to align their actions with those of their neighbors. The interaction network defines both the intensity and direction of coordination motives. This framework captures the key trade-offs in the motivating examples: firms adjusting prices to aggregate costs or demand while aligning with specific partners, or dealers setting quotes to track asset values while maintaining spread relative to counterparties.
The analysis reveals a stark non-monotonicity between the value of public information and network connectivity. While public information is beneficial when interactions are uniformly weak or uniformly strong, it can be detrimental in the intermediate case where interactions are concentrated among a few hub agents. I show that providing noisy public information harms welfare when the network exhibits sharp structural heterogeneity, specifically in core-periphery networks characterized by small, dense cores. This result arises from a core-driven distortion. Core agents, prioritizing internal coordination, anchor heavily on the public signal and deviate from the fundamental state. This deviation exacerbates the strategic trade-off for peripheral agents: as the core’s actions and the fundamental state diverge, peripheral agents are induced to sacrifice alignment with the fundamental to maintain coordination with the core. When the core is small and dense, the welfare loss arising from this intensified trade-off dominates the direct informational benefits of the signal.
These topological distinctions map directly to real-world market structures. While hub-and-spoke architectures with sparse inter-hub connectivity characterize production networks (e.g., Atalay et al., 2011; Mizuno et al., 2014; Bernard et al., 2019), tiered architectures with dense cores are typical of OTC markets (e.g., Craig and Von Peter, 2014; Hollifield et al., 2017; Li and Schürhoff, 2019). Since the welfare impact of public disclosure depends on core density, a regulator operating without precise knowledge of the underlying network risks implementing counterproductive policies. My findings thus underscore the potential for substantial welfare gains from acquiring granular network data.
More formally, I formulate the problem within a linear-quadratic-Gaussian framework, which admits a unique linear equilibrium. In this equilibrium, an agent’s responsiveness to public information is proportional to her Katz-Bonacich centrality. This centrality measure aggregates an agent’s total connectivity by summing all discounted paths of any length emanating from her. Intuitively, these paths represent the agent’s indirect coordination motives. For instance, a path to agent $k$ through neighbor $j$ captures $i$’s derived incentive to align with $k$ to coordinate with $j$. These indirect motives amplify the equilibrium weight on the public signal because it serves as the common focal point for such higher-order alignments. Thus, agents with higher centrality place greater weight on the public signal.
The first main result establishes that providing noisy public information reduces an agent’s equilibrium payoff if and only if her Katz-Bonacich centrality is sufficiently low relative to the aggregate centrality of her neighbors. To illustrate, consider a simple three-agent network comprising Ann, Bob, and Carol. Bob and Carol are reciprocally connected, forming a dense core, while Ann connects solely to Bob. Reflecting his high centrality, Bob excessively responds to the public signal, rendering his action noisy relative to the fundamental. Ann, in her peripheral position, is forced into a strategic trade-off: to maintain coordination with Bob, she must absorb this induced volatility into her own payoff. When the centrality gap between Ann and Bob is large, the cost of this imported volatility outweighs the direct informational benefit of the signal, resulting in a net payoff loss for Ann.
These distributional effects translate into a precise condition for an aggregate welfare loss. I show that the provision of noisy public information becomes socially detrimental if and only if the network features agents who have a combination of disproportionally high Katz-Bonacich centrality and high in-degree. This condition decomposes the mechanism of the welfare loss: high centrality determines the magnitude of the distortion (i.e., how much the agent over-reacts to public information), while high in-degree determines the scale of its transmission (i.e., how many others intend to follow her action). High centrality alone is insufficient to generate aggregate loss; the distorted agent must also be influential enough to propagate this volatility to a broad set of neighbors.
Building upon this insight, I also identify a novel source of inefficiency in endogenous information sharing. In an extension allowing for voluntary disclosure, I show that agents may rationally withhold information to avoid this core-driven distortion. A peripheral agent, foreseeing that releasing her signal will induce highly central neighbors to anchor aggressively on it, may choose to suppress the signal to avoid importing the resulting volatility. This equilibrium is often socially inefficient because the agent fails to internalize the coordination benefits the signal would provide to the core, leading to an under-provision of public information.
Finally, while the finding that public information can reduce welfare parallels the seminal result of Morris and Shin (2002), the underlying economic mechanism is distinct. Morris and Shin (2002) rely on a preference misalignment where the social planner cares solely about alignment with the fundamental. In contrast, the present result assumes a utilitarian planner, and the welfare loss arises only when there is significant structural heterogeneity in the interaction network, which is absent in their symmetric framework.
@@ 2 Literature
The literature on the social value of information, pioneered by Morris and Shin (2002), establishes that public information can reduce welfare when private and social incentives diverge (see Pavan and Vives (2015) and Angeletos and Lian (2016) for overviews). Analyzing general symmetric linear-quadratic games, Angeletos and Pavan (2007) formalize this insight, attributing the value of information to the discrepancy between equilibrium and socially optimal degrees of coordination. Ui and Yoshizawa (2015) subsequently provides a full characterization of the value of information for this class of games. Extensions have identified specific frictions that complicate this baseline, such as dynamic learning externalities (Amador and Weill, 2010) or the crowding out of private information acquisition (Colombo et al., 2014).
Despite these advances, the existing literature predominantly assumes ex-ante payoff symmetry. [^1] Exceptions include Huo and Pedroni (2020) on dynamic games and Miyashita and Ui (2025) on information design, though their focus differs from the analysis here. I depart from this benchmark by analyzing a general framework of heterogeneous interaction structures. My analysis reveals that such heterogeneity creates a pronounced structural wedge between private and social incentives, substantial enough to reverse the sign of the welfare effect even under a utilitarian planner. This implies that the key determinants of the value of information lie not in global preference parameters, but in the granular network structures through which agents interact.
This paper also relates to the literature on endogenous information acquisition and sharing. The seminal models of endogenous information acquisition in beauty contest games identify strategic complementarities: the marginal value of a signal increases with the number of other agents who also observe it (Hellwig and Veldkamp, 2009; Myatt and Wallace, 2012). Recent contributions, such as Myatt and Wallace (2019) and Leister (2020), extend this framework to network games built upon Ballester et al. (2006). In a related model, Herskovic and Ramos (2020) demonstrate that coordination motives drive agents to endogenously form core-periphery networks to facilitate information exchange. These studies generally find that connectivity reinforces the incentive for shared information, preserving the intuition that agents prefer to coordinate on common signals.
In contrast, I show that sharp core-periphery structures reverse this logic. In such networks, the negative externality of induced aggregate volatility dominates the incentive to coordinate on common signals. This identifies a novel source of inefficiency: peripheral agents may rationally withhold signals to avoid the volatility caused by public disclosure.
More broadly, this paper contributes to the growing literature on optimal interventions in network games. Ballester et al. (2006) focus on structural interventions, identifying the “key player” whose removal maximally impacts aggregate activity. Galeotti et al. (2020) study incentive interventions, characterizing optimal adjustments to agents’ marginal returns. In the domain of information design, Miyashita and Ui (2025) develop a semidefinite programming approach to characterize the optimal information structure in general linear-quadratic-Gaussian games. Unlike these approaches, which typically derive agent-specific optima, I focus on public disclosure, which is the standard policy instrument in decentralized markets where fine-tuned intervention is often infeasible.
@@ 3 Model
I introduce a beauty contest game in which agents’ coordination motives are represented as a directed and weighted network.
@@
Information Structure. The information structure is standard, following Morris and Shin (2002). The payoff state $\theta$ follows an improper uniform prior distribution over the real line. Each agent $i$ observes a signal vector $\mathbf{s}_{i}=(x_{i},y)$, where $x_{i}$ is a private signal observed only by $i$, and $y$ is a public signal observed by all agents. Specifically, I define $x_{i}\equiv\theta+\varepsilon_{i}$ and $y\equiv\theta+\varepsilon_{y},$ where $\varepsilon_{i}$ and $\varepsilon_{y}$ are normally distributed with
$\mathbb{E}[\varepsilon_{i}]=\mathbb{E}[\varepsilon_{y}]=0,\mathrm{Var}[\varepsilon_{i}]=\tau_{x}^{-1},\mathrm{Var}[\varepsilon_{y}]=\tau_{y}^{-1},$
and the noise terms $\varepsilon_{i},\varepsilon_{y}$ and $\theta$ are all jointly independent. The relative precision of private signals is denoted by $\gamma\equiv\tau_{x}/(\tau_{x}+\tau_{y})\in(0,1)$.
@@
where $\mathbb{E}_{i}[\cdot]\equiv\mathbb{E}[\cdot\mid\mathbf{s}_{i}]$ is the conditional expectation operator, conditional on $i$’s information $\mathbf{s}_{i}$. Consider linear strategies of the form $a_{i}=b_{i}y+(1-b_{i})x_{i}$ for all $i\in N$. [^2] Focusing on this form of linear strategies is without loss of generality. In Section A.1, I show that the equilibrium slopes must sum to one for all agents in any linear equilibrium. Plugging these linear strategies into (2) yields
@@
where $I$ is an $n\times n$ identity matrix, and $\mathbf{b}=(b_{1},\ldots,b_{n})^{\prime}$ and $\mathbf{1}=(1,\ldots,1)^{\prime}$ are $n$-dimensional column vectors. The system of linear equations (5) has a unique nonnegative solution for $\mathbf{b}$ because the inverse matrix $(I-\gamma G)^{-1}$ exists and is nonnegative. [^3] By the Perron-Frobenius theorem, $G$ has the largest nonnegative real eigenvalue $\rho(G)$ and $\rho(G)\leq\max_{i\in N}\sum_{j\neq i}g_{ij}<1$ by assumption. Also, by Theorem III of Debreu and Herstein (1953), $(I-\gamma G)^{-1}$ exists and is nonnegative if and only if $\gamma\rho(G)<1$. Combining these observations proves the claim. The vector $\mathbf{c}(\gamma,G)\equiv(I-\gamma G)^{-1}\mathbf{1}$ is called the Katz-Bonacich centrality vector in the network literature.
@@
In equilibrium, agent $i$’s direct and indirect coordination motives are positively associated with her responsiveness to the public signal. Intuitively, when $i$’s coordination motive strengthens, the commonality of the public signal becomes more valuable for aligning her action with her neighbors’ actions. Moreover, if $i$’s neighbors have stronger coordination motives, they will respond more to the public signal, further increasing the value of the public signal as a coordination device. Thus, not only an agent’s own coordination motive but also those of her neighbors (and neighbors’ neighbors, etc.) positively affect her responsiveness to the public signal. [^4] The relative precision of private information, $\gamma$, acts as the equilibrium discount factor in the Katz-Bonacich centrality. This role arises from how different orders of agents’ expectations about $\theta$ (e.g., $\mathbb{E}_{i}[\theta]$ and $\mathbb{E}_{i}[\mathbb{E}_{j}[\theta]])$ respond to changes in $\gamma$. A more detailed intuition is provided in Murayama (2025). Following the standard practice in this literature, I focus on this unique linear equilibrium in the analysis.
@@ 4 Results
The main results show that public information provision can be detrimental to an individual agent and aggregate welfare, depending on the geometry of interaction networks. Most proofs and derivations are provided in the appendix.
@@ 4.1 Individual Payoff
Let $\Delta U_{i}(\tau_{x},\tau_{y},G)\equiv U_{i}(\tau_{x},\tau_{y},G)-U_{i}(\tau_{x},0,G)$ be the effect of public information provision on agent ii. [^6] Abusing notation, I write $U_{i}(\tau_{x},0,G)=-\tau_{x}^{-1}(1+d_{i}^{out})$ for agent $i$’s equilibrium payoff in the absence of public information. The following result is immediate from (11).
@@
The detrimental effect on Ann, stemming from the increased commonality of a signal between Bob and Carol, is reminiscent of a mechanism highlighted by Morris and Shin (2002). Morris and Shin show that when agents’ common coordination motive (analogous to $\beta$ in my model) is strong enough and the relative precision of private information is sufficiently high, public information provision increases the action-state distance, i.e., $\mathbb{E}[(a^{*}_{i}-\theta)^{2}]$. Similarly, the increased commonality of information between Bob and Carol can make Bob’s equilibrium action distant from $\theta$. [^8] Intuitively, this occurs because Bob and Carol’s coordination motive is so large that they choose to be much more responsive to the public signal than they would be when their objectives are to align their actions only with $\theta$. This is not desirable for Ann, whose objective is to simultaneously align her action with $\theta$ and Bob’s action.